OpenStudy (anonymous):

determine if the equation has the same slope as the equation. y=2x-4. 1. y=2x+4 2. -2x+3 3. y= 4x-2 4. y=3x-4 and please explain how you did one! thankyou!

OpenStudy (anonymous):

1. same slope 2. not the same slope 3. not the same slope 4. not the same slope

OpenStudy (anonymous):

for y = mx + b, m is the slope so for y=2x-4, slope m = 2 Therefore 1. y=2x+4 same slope (m= 2) 2. -2x+3 not same slope (m= -2) 3. y= 4x-2 not same slope (m= 4) 4. y=3x-4 not same slope (m= 3)

OpenStudy (anonymous):

oh thankyou that was a big help i understand it so much better now!

OpenStudy (anonymous):

u r welcome☺☻